Calculate the no. of exchange pairs of electrons present in configuration of Cu acc to Auf-Bau Principle considering 3d and 4s.

Grade:11

1 Answers

Suraj Prasad IIT Patna
9 years ago
The crystal structures of the dimer form of copper(II) tetraphenylporphyrin N-oxide, [Cu(tpp-N-O)]2 (3-dimer), and zinc(II) tetraphenylporphyrin N-oxide, [Zn(tpp-N-O)]2 (4-dimer), were established. The geometry at the copper ion in 3-dimer is essentially square-pyramidal with one oxygen bridge occupying the apical site, giving a much larger Cu-O bond distance compared to those at the basal plane. The respective Cu...Cu distance and Cu-O-Cu angle in the core of 3-dimer are 3.987(4) A and 148.1(3) degrees. The Zn(1) atom in 4-dimer has a distorted square-pyramidal [4 + 1] coordination geometry that gives a tau-value of 0.19. The respective Zn...Zn distance and Zn-O-Zn angle in the dimeric unit of 4-dimer are 4.025(3) A and 148.1(2) degrees. The 3-dimer displays axial X-band electron paramagnetic resonance spectral features (Es = 0) in the powder state at 4 K, giving g parallel = 2.51 (A(parallel,s) = (9.6 +/- 0.2) x 10-3 cm(-1)) and g(perpendicular) = 2.11 and in the same powder state at 293 K giving Ds = 0.0731 cm(-1) (as derived from DeltaMs = 1 lines) or 0.0743 cm(-1) (as derived from the DeltaMs = 2 lines). In addition, 3-dimer displays a DeltaMs = 2 transition at g = 4.17 indicating the presence of spin-exchange coupling. The anisotropic exchange interaction (Ds(ex)= 0.132 cm(-1)) gives the main contribution to Ds in 3-dimer. The theoretical fit of the susceptibility and effective magnetic moment data of 3-dimer in the temperature range of 5-300 K gives 2J = 68 cm(-1), g = 2.01, p = 0.06, and a temperature-independent paramagnetism of 10(-6) cm3 mol(-1). This magnetic susceptibility data indicates that the copper(II) ions in 3-dimer are coupled in a ferromagnetic manner with the ground-spin triplet stabilized by 68 cm(-1) with regard to the singlet.